Some Gun fanciers made their own bullet which involves melting and casting the lead slugs. How much heat transfer is needed to raise the temperature and melt 0.5kg of lead starting from 25°C?
1
Expert's answer
2019-07-29T08:18:31-0400
"Q=m(L+c\Delta T)"

"Q=0.5(24500+(128)(327.3-25))=31600\ J=31.6\ kJ"